What is the value in year 3 of a $600 cash flow made in year 6 if interest rates are 8 percent?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

600=P*(1.08)^3

P=600/1.08^3

which is equal to

=$476.30(Approx).